### v1

# LayerCollapse: Adaptive compression of neural networks

Abstract
Handling the ever-increasing scale of contemporary deep learning and transformer-based models poses a significant challenge. Although great strides have been made in optimizing model compression techniques such as model architecture search and knowledge distillation, the availability of data and computational resources remains a considerable hurdle for these optimizations. This paper introduces LayerCollapse, a novel alternative adaptive model compression methodology. LayerCollapse works by eliminating non-linearities within the network and collapsing two consecutive fully connected layers into a single linear transformation. This approach simultaneously reduces both the number of layers and the parameter count, thereby enhancing model efficiency. We also introduce a compression aware regularizer, which compresses the model in alignment with the dataset quality and model expressiveness, consequently reducing overfitting across tasks. Our results demonstrate LayerCollapse’s effective compression and regularization capabilities in multiple fine-grained classification benchmarks, achieving up to 74% post training compression with minimal accuracy loss. We compare this method with knowledge distillation on the same target network, showcasing a five-fold increase in computational efficiency and 8% improvement in overall accuracy on the ImageNet dataset.
1 Introduction
The expanding scope of Deep Learning (DL) applications has led to an escalating demand for more accurate models, resulting in an ongoing rise in their complexity. This surge in complexity, however, presents a challenge: it increases the execution costs, making it difficult to deploy these models in real-world scenarios on standard hardware. For example, transformer-based deep models while improving accuracy, present new computational challenges. Fortunately, modern trained neural networks are shown to include excess redundancies that can be removed without losing accuracy. This potential for reduction has catalyzed research in model compression strategies such as pruning [24][19], quantization [15][18][29][20], and model distillation[14], focusing on the effective detection and elimination of these superfluous components. Despite the amazing progress in the field of model compression, these optimizations demand the availability of data and high computational resources which further hinders their practicability.
Multi-Layer Perceptrons (MLPs) [11] is a key architectural component of modern deep neural networks. MLPs are regarded by many as the first contribution to deep learning, first introduced in 1965 by Ivakhnenko [17][32]. In its simplest form, a $K$-layer MLP can be described as $K$ linear transformations separated by some non-linear operations. Equivalently, an MLP can also be described as a sequence of fully connected (a.k.a., dense), layers separated by non-linear layers. Often, MLP layers serve as the final decision-making components, such as classification heads. MLPs are central to the success of deep learning models in diverse architectures including Convolutional Neural Networks [33], Vision Transformers [7], and MLP-Mixers [35].
Because of the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to the MLP layers. For instance, MLPs constitute over 60% of the model’s total parameters for the popular models discussed in [7][35][33].
Table 1: Percentage of model parameters and MACs reductions attainable by collapsing their MLP layers
Model % Params % MACs
ViT T/16 $54.3$ $56.7$
ViT B/16 $57.3$ $57.9$
ViT L/16 $57.9$ $58.2$
Mixer B/16 $85.0$ $85.2$
Mixer L/16 $86.5$ $86.5$
VGG 16 $71.2$ $0.6$
MLPs contain non-linear activations, e.g., ReLU layers, which separate sequences of linear transformations. If the non-linear layers were absent, the consecutive linear layers could be condensed into a single layer.To showcase the immense potential of computation and storage reductions when linearizing MLPs, we calculate this reduction in the architectures presented in [7][35][33]. Table 1 demonstrates the reduction percentage in the number of parameters and MAC operations when we replace the MLP’s with a linear transform in relevant vision architectures.
The core thesis of our work revolves around non-linear activations. We argue that the expressiveness of an MLP is primarily governed by its non-linear components. By modulating these non-linear characteristics, we can orchestrate a continuum of model expressiveness, ranging from highly complex to near-linear transformations. Figure 1 demonstrates this idea using a simple regression task and an MLP with 2 layers. As illustrated, having complex models with lack of a sizeable dataset will lead to overfitting. By modulating the non-linear activation layer one can regulate model’s expressiveness.
Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting.
Classically, regularization strategies such as the $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, ensuring robust generalization and preventing overfitting. These techniques also contribute to compression-aware training. In similar veins, Lemaire et. al. describe a budget aware regularizer [23] and Nguyen et. al. introduce a regularizer for quantization aware training [29]. Extending this principle, we introduce a compression aware regularizer by using the nature of ReLU activations. This regularizer works by shifting the slope of Relu towards linearity. This not only facilitates the fusion of successive linear layers, conserving computational resources and storage, but also can lead to better model generalizability.
This paper presents several key contributions to the field of neural network compression and regularization:
  • Introduction of a novel post-training technique, called LayerCollapse, that combines two fully connected layers into one, reducing model size and complexity.
  • Development of a novel compression aware regularization loss with the objective of promoting shallower structures during training.
  • Proof-of-concept experimentation on a diverse set of datasets benchmarks, in vision and NLP domaains, showing post training compression up to 74%. Furthermore we show improvement in generalizability when training with proposed compression aware regularizer while decreasing size upto 50%. Lastly we compare the computational performance of knowledge distillation and LayerCollapse showing an 8% final result improvement while using 80% less computational resources.
  • Providing an open-source library that enables easy implementation of LayerCollapse, compatible with PyTorch. [^1] https://github.com/sohail2810/LayerCollapse
2 Related work
2.1 Compression
In the realm of deep learning, compression techniques aim to reduce computational demands while ensuring model precision. With the emergence of foundational models and the tapering of Moore’s law, the correct utilization of resources is a paramount concern. Predominantly, the methods employed for model compression are pruning [10], quantization [15][18], and distillation [14], each with its unique approach toward model optimization.
Pruning eliminates neurons whose impact on model performance are marginal or negligible. A big body of work explores diverse strategies to discern which neurons are expendable, including magnitude [24] and gradient [26][41] based methods. Another body of work looks into pruning aware training (PAT). In these methods a regularization term ensures structural sparsity when training [23][19][38][1]. This reduces the loss of accuracy when pruning the model post training.
Quantization constrains the parameter range and maps it into integers, thereby reducing the model’s computational and storage requirements. This methodology has been explored and implemented in various studies, providing numerous approaches toward effective quantization [9][29]. To minimize the loss induced by quantization it is standard practice to do quantization aware training (QAT). In short, we regularize the model weights during training to stay within a a close range of their quantized value [29].
Distillation[14], within another compression category, initially seeks an optimized architecture via Neural Architecture Search (NAS), a procedure notably resource-intensive in itself. The target model is trained using this new architecture, utilizing the dataset and the original model as a teacher. This process is sometimes called the student-teacher paradigm. It’s important to note that both NAS and distillation are resource-draining processes and require access to training data, which may not always be accessible.
In section 4, we compare LayerCollapse against distillation, highlighting its respective merits and applicability.
2.2 Regularization
In the field of deep learning, achieving the right balance between underfitting and overfitting is a pivotal challenge. Overfitting occurs when a model excessively learns from the noise in the training data, reducing its ability to generalize to new datasets. This issue is typically managed through the bias-variance trade-off, a principle that aids in creating models capable of effective generalization. A model with high variance is overly sensitive to small fluctuations in data, leading to unstable predictions. This trade-off highlights the necessity of maintaining an equilibrium between model complexity and the volume and integrity of the data. To counteract overfitting and enhance performance on test data, it’s crucial to implement explicit regularization strategies. These methods ensure the model’s robustness and consistent performance across both training and test datasets. Research by Yang et al. demonstrated that $\ell_{1}$ regularization can serve as a pruning-aware training method to sparsify the model, thereby boosting its pruning efficiency[40]. Ding et al. employed a specialized $\ell_{2}$ regularization to selectively eliminate less significant channels during training, aiming for subsequent pruning [4]. Additionally, Nguyen et al. introduced a regularization technique specifically for quantization-aware training [29]. Drawing inspiration from Santos et al.’s classification framework, we categorize model generalization techniques into three distinct categories: “Ensembling”, “Explicit Model Regularizers”, and “Dataset Design”[31].
Ensembling techniques often revolve around the integration of multiple models or sub-models to enhance performance and generalization. Within this domain, methods such as Dropout and its variants [34][5][8] play a pivotal role.
Dataset Design centers on refining the input data and labels to achieve model generalization. This includes data augmentation techniques such as transformative actions (random cropping, blurring, and rotation) seen in works like Bag of Tricks [13], FixRes [36], Cutout [3], CutMix [42], and AutoAugment [25].
Explicit regularizers modulate model configurations either through architectural adjustments or by penalizing certain weights to prevent overfitting. These methods follow the principle of Occam’s razor. They assume a prior knowledge that simpler models are more likely to be robust. This category encompasses classical regularization techniques like L0, L1, and L2 (weight decay) [27]. Our proposed method, falls into the same category. In section 3, we demonstrate how we use this principle in designing our regularization term. While classical norms influence weight magnitudes to mitigate overfitting, LayerCollapse uniquely stands out in its ability to adjust the model’s depth and structure.
3 Methodology
Figure 2: Overview of the LayerCollapse process: $(a)$ The compression aware the regularization loss shifts the PReLU towards linearity. $(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
This section introduces LayerCollapse, an innovative compression method that adaptively reduces the model’s depth through regularization. This methodology aims to optimize the size of the model based on the quality of the dataset, resulting in enhanced generalization. Our proposed framework seamlessly integrates regularization and compression.
Unlike pruning [24] and quantization [15] which require the compressed model to be finetuned, LayerCollapse requires the finetuning prior to the compression phase.
As depicted in Figure 2, during a compression aware training or finetuning, we regularize a layer’s complexity by linearizing its activation function. We define Collapsing an MLP on a model that has shifted at least one of its non-linear layers into linearity. As depicted in Figure 2 Linearizing a non-linear activation within a model allows for the collapse of the two surrounding linear layers into one linear layer, effectively reducing the depth. The resulting model has fewer parameters and a smaller depth, requiring less resources for execution. For post training compression, one needs to run a few cycles of finetuning using the LayerCollapse regularizer to ensure the linearization loss is marginal.
We define a Collapsible MLP to be two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) [12] in the middle. The Parametric Rectified Linear Unit (PReLU) can be expressed as:
$PReLU_{\alpha}(x)=\max(0,x)+\alpha\min(0,x)$ (1)
where $\alpha$ represents the slope of the PReLU function when x is negative. Output of a basic two-layer collapsible MLP is described as:
$Y_{\alpha}=W_{2}(PReLU_{\alpha}(W_{1}X+b_{1}))+b_{2}$ (2)
where, $W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively. $\alpha$ can be used to make the activation linear. We define $Y_{linear}$ as:
$Y_{linear}=W_{2}W_{1}X+W_{2}b_{1}+b_{2}$ (3)
We define the error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta>0$
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (4)
where,
$x^{\delta}\equiv\inf\left\{x\in\mathcal{R}^{n}\mid P\left\{|X|>|x|\right\}<\delta\right\}$ (5)
$C=\sigma_{max}(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2}$
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, error is bounded proportional to $(1-\alpha)^{2}$ with probability one. We capitalize on this by incorporating $(1-\alpha)^{2}$ into our loss function. This loss term serves to align $Y_{\alpha}$ with $Y_{linear}$, compressing the model into a shallower model as demonstrated in Figure 2. In simpler terms, when $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ with an error bounded by a constant multiplied by $(1-\alpha)^{2}$.
In practical applications, Dropout [34] and BatchNormalization [16] are frequently used between the layers of a network. Dropout works by temporarily deactivating random neurons during training, which helps spread the learning across different neurons to avoid over-reliance on any particular subset of them. It is important to note that Dropout is not active during model deployment, meaning it behaves as if it is not there. This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out when considering it.
BatchNormalization is a process described by the formula:
$Y=\frac{x-E[X]}{\sqrt{Var[X]}}\times\gamma+\beta$ (6)
where, $\gamma$ and $\beta$ are learnable parameters, the $E[X]$ and $Var[X]$ are the mean and variance of the current batch, which we sometimes adjust with a method called Polyak averaging to keep the batch-to-batch changes smooth. Algorithm 1 outlines the specific steps to be taken for adjusting weights and biases when BatchNormalization is present.
To measure how well our compression method works, we use the concept of “compression gain”. This is calculated as the percentage reduction in the number of parameters of a layer, excluding bias and BatchNormalization parameters. The formula for compression gain is:
$Gain=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)
where, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of a 2-layer MLP, respectively.
An interesting aspect of this formula is that the gain can sometimes be negative. Positive gain happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum. This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see much benefit from our layer collapsing method. On the other hand, most modern MLP designs, which tend to expand the layer size (“widening” structure), are likely to experience significant positive gains from our method.
3.1 Adaptation for Convolutional Layers
Convolutional layers play a pivotal role in numerous neural network architectures. Essentially convolution can be formulated as sparse fully connected layers with some redundancies. Their nature as linear transforms permits representation via circulant matrices. A circulant matrix is a matrix where the row vectors have the same elements and each row vector is rotated one element to the right relative to the preceding row vector. An interesting property of these matrices is that their product also forms a circulant matrix, resulting in another convolution layer. The kernel size of the resulting convolution layer, denoted as $k$, is calculated by $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers. The compression gain can be calculated by:
$Gain_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
where, $c_{in}$, $c_{hidden}$, and $k_{1}$ correspond to the input channels, output channels, and the kernel size of the initial convolution layer, respectively. Meanwhile, $c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer. However, an essential note when applying LayerCollapse to common CNN architectures, such as ResNet, is that the gain often turns out to be negative. To achieve a positive gain, it is imperative for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$. This observation highlights the need for careful configuration when attempting to collapse layers in such architectures. Overall, this framework offers a principled approach to explore the potential benefits of LayerCollapse on convolution layers, illuminating the optimization potential for convolutional neural networks.
3.2 Prior Distribution Regularization
We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework. Aligning with Occam’s razor, which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$. This setting effectively limits the two layers to a linear space, reducing their functional complexity, as illustrated in Figure 1. This choice is backed by empirical results given in section 4. MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function. This is referred to as the prior loss or regularization loss. Following this principal the added regularization term is:
$\mathcal{L}_{\text{reg}}=lc\times(1-\alpha)^{2},\text{}$ (9)
where $lc$ is the regularization constant.
Algorithm 1 Layer Collapse
1:function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$)
2:     if $\alpha\leq\tau$ then
3:         $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
4:         $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
5:         $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
6:         return $W_{\text{new}},b_{\text{new}}$
7:     else
8:         return un-collapsible
9:     end if
10:end function
4 Experiments
In this section we evaluate our method on both post training compression, and compression aware training with our regularization. Initially, we explore the standalone capability of LayerCollapse in post training compression of streamlined vision models including vision transformer, mlp-mixer and VGG. We then compare LayerCollapse with Knowledge Distillation on the cost of training and achievable results on a target network. Subsequently, we demonstrate the effectiveness of compression aware training in an ablation study on mid-sized datasets: TinyImageNet [22], and CIFAR100 [21]. Further extending the scope of our investigation, we exhibit preliminary results on the BERT [2] language model to provide a proof-of-concept for the generalizability of LayerCollapse in domains beyond computer vision, such as natural language processing (NLP). These experiments aim to demonstrate LayerCollapse’s versatility and its potential to enhance efficiency and performance in various machine learning contexts.
Model Collapsed Layers No. of Params (M) Top1 Acc. (%) Top5 Acc. (%) MACs (G)
ViT T/16 0 5.72 75.35 93 1.08
1 5.46 75.79 93 1.02
2 5.20 75.35 92 0.97
3 4.94 74.42 92 0.91
ViT S/16 0 22.05 81.38 96 4.24
1 21.02 80.96 96 4.04
2 19.98 79.74 95 3.84
ViT B/16 0 86.57 81.43 96 16.85
1 82.44 82.17 96 16.04
2 78.3 80.31 95 15.23
ViT L/16 0 304.33 84.80 98 59.66
2 289.64 84.22 97 56.77
Mixer B/16 0 59.88 76.47 92 12.61
2 55.64 76.11 92 11.71
4 51.39 75.67 91 10.81
VGG19 0 143.67 74.21 92 19.65
2 45.11 71.22 88 19.55
VGG16 0 138.36 73.36 92 15.48
2 39.80 67.16 88 15.39
VGG13 0 133.05 71.58 90 11.32
2 34.49 67.21 88 11.22
VGG11 0 132.86 70.37 90 7.62
2 34.31 66.36 88 7.52
Table 2: Post training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
4.1 Compression
Our compression experiments are performed on ImageNet-1k classification task [30]. Unlike pruning [24] and quantization [15] compression methods which require finetuning on the compressed method to recover accuracy, LayerCollapse utilizes finetuning with the regularizer to ensure accuracy. Afterwards the compression phase will endure no accuracy loss. We first assess the stand-alone compression efficacy of LayerCollapse. This is followed by a comparative analysis of LayerCollapse and Knowledge Distillation, focusing on training efficiency and resource utilization.
4.1.1 Stand-Alone Compression
Dataset. We employ the ImageNet-1k [30] dataset for both fine-tuning and validation processes. It comprises over 1.2 million images for training and 50,000 for validation.
Implementation. We adapt pre-trained models from the TIMM library [39] for MLP-Mixer [35] and ViT [7], alongside VGG variants from torchvision [28]. ViT and Mixer architectures use GeLU activations in their original implementation. In target layers, these are converted to ReLU to align with the requirements of LayerCollapse. The fine-tuning follows a consistent procedure across models, using Stochastic Gradient Descent (SGD) with momentum and a learning rate of $5e^{-4}$ as the optimizer. The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and uncollapsed models output and the LayerCollapse regularization term (9) with $lc=0.2$. This process is capped at 20 epochs. Each layer for all ViT experiments use 5 epochs. Number of epochs for VGG architectures are given in Table 3. The MAC operations are all calculated by propagating a random $224\times 224$ image through the model. Further details are enumerated in Appendix B.
Sensitivity Analysis. To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 3. This removal of layers provides insight into the impact of incremental compression on the model’s accuracy. Interestingly, we observe an initial increase in accuracy upon the collapse of the first layer, indicating an improvement in generalization compared to the original model.
Figure 3: Layer-wise collapse accuracy and parameter reduction analysis for ViT-T/16.
Model Method Top1 Acc(%) Top5 Acc(%) No. of Parameters (M) Cost Epochs
Initial Final
vgg19 LC 71.22 88 143.67 45.11 11h 5
KD 63.42 85 51h 23
vgg16 LC 70.01 89 138.36 39.80 14h 7
KD 59.63 82 31h 16
vgg13 LC 67.21 88 133.05 34.49 16h 10
KD 58.26 81 45h 28
vgg11 LC 66.36 88 132.86 34.31 20h 20
KD 58.64 81 42h 42
Table 3: A comparative analysis of Layer Collapse (LC) and Knowledge Distillation (KD) methods applied to VGG models.
Discussion. Table 2 illustrates that LayerCollapse achieves a substantial reduction in model size—up to 74% in VGG11—and computational demands with minimal accuracy degradation. Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method. ViT models, the T/16 and B/16 variants, get an increase in accuracy with the first layer collapse. This can be attributed to an increase of generalization in simpler models. In ViT T/16, S/16, and B/16 we get roughly 5 percent reduction in model size per LayerCollapse. In all ViT variants we get a consistent reduction of over 10% in size and complexity with less than a 1% accuracy loss. In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist. MLP-Channel has a significantly bigger size. We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort. In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected. In the Mixer architecture we get 14% reduction in size with less than a 1% accuracy degradation. Notice that the Mixer architecture is the only one that gets a linear reduction in MAC operations with each pair collapsed. We used the idea first introduced in Knowledge Distillation to enhance the performance of our finetuning procedure. Using the KL-Divergence loss, we speedup the gradient flow and increase the convergence speed.
4.2 LayerCollapse vs Knowledge Distillation
Setup. Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository [28], with ImageNet-1K [30] providing a substantial dataset for training and validation. These variants are chosen for their prevalent use in compression studies and their capability to learn without prior training. To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods. The rest of the training is set up identical to section 4.1.1 for both LayerCollapse and Knowledge distillation.
Comparison. The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set. The computational costs are quantified on an Nvidia RTX A6000 GPU.
Discussion. Referencing Table 3, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation. This trend is more pronounced in larger models, which exhibit greater compression potential, and result in up to 15% accuracy gain with 50% of tuning cost. It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical. LayerCollapse emerges as a unique method for model compression through architectural modification, offering significant cost advantages as demonstrated in Table 3. However, it is important to recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures.
4.3 Ablation Study of Regularization
We provide an in-depth analysis of how our regularization approach impacts both the generalization and compression abilities in training image classifiers from scratch.
Dataset. Our study utilizes the CIFAR100 [10] and TinyImageNet [22] datasets, which are mid-sized and rich in diversity. The CIFAR100 dataset includes 60,000 images across 100 classes, and TinyImageNet, a subset of the larger ImageNet, comprises 100,000 images across 200 classes.
Implementation. The experiments are conducted using the VGG architecture because of its effective learning capabilities without pre-training and competitive performance in benchmarks. For consistency, all experiments utilize the same setup and hyper parameters. Note that the goal of this study is not to show state of the art performance using our proposed regularization, but to show its effectiveness in reducing over-fitting and final model size. In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropout[5], and batch normalization[16], have been used. This is to show the added contribution of 9. Training is done using Stochastic Gradient Descent (SGD) with momentum a learning rate scheduler, trained for a total of 400 epochs per experiment. The regularization is applied to 40% of the layers, starting with the latest layers, with the strength of 0.05 across all experiments. extensive experiment details can be found in Appendix B.
Discussion. The results, documented in Table 4, demonstrate the benefits of incorporating compression aware regularization. We observe an enhancement in top-1 accuracy by as much as 2% with 50% reduction in model size. The learning curve presented in Figure 4 suggests a narrowing gap between training and validation accuracy, indicating a reduction in overfitting and highlighting the effectiveness of our regularization strategy in improving generalization.
4.4 LayerCollapse in NLP
We extended LayerCollapse to the NLP domain. This section presents our initial findings using the BERT-Base model [2], a well-established pre-trained NLP model.
Setup In our evaluation we use the General Language Understanding Evaluation on the Microsoft Research Paraphrase Corpus (GLUE-MRPC)[37][6]. This benchmark assesses a model’s capacity to discern paraphrased sentences, determining if they are equivalent. Further details are found in Appendix B
Discussion We report our results in Table 5. These include accuracy metrics and the number of parameters for each layer of BERT-Base modified by LayerCollapse. The experiment shows a promising 19% reduction in parameters with a minimal accuracy loss of less than 3%. Notably, we exclusively use the data in the GLUE-MRPC task in the LayerCollapse procedure. These preliminary results in the NLP sector suggest the versatility and potential of LayerCollapse across different domains, not limited to computer vision. They point to the method’s adaptability and effectiveness in optimizing model parameters while preserving accuracy without the need for pre-train dataset.
Figure 4: Evaluating the impact of LayerCollapse regularization on CIFAR100, we compare the top-1 accuracy of VGG11 both with and without this regularization technique.
Table 4: Number of parameters after Collapse and Top-1 Accuracy of various models with and without LC on CIFAR-100 and Tiny ImageNet.
Model CIFAR-100 Tiny ImageNet
#Param Top-1(%) #Param Top-1(%)
VGG11 $11.7$M $59.7$ $18.4$M $60.2$
VGG11(LC) $9.3$M $63.5$ $9.6$M $61.7$
VGG13 $11.9$M $66.1$ $18.6$M $50.3$
VGG13(LC) $9.5$M $67.8$ $9.8$M $52.2$
VGG16 $17.2$M $63.3$ $23.9$M $48.9$
VGG16(LC) $14.8$M $64.1$ $15.1$M $49.7$
VGG19 $22.5$M $62.5$ $29.6$M $46.8$
VGG19(LC) $20.1$M $63.8$ $20.4$M $48.2$
Table 5: Layer-wise analysis of accuracy and parameters for Bert-Base via LC
Layers Accuracy (%) Params (M)
0 $85.0$ $109.40$
1 $85.8$ $105.35$
2 $84.1$ $101.20$
3 $84.3$ $97.10$
4 $82.1$ $92.95$
5 $82.3$ $88.80$
6 $82.3$ $84.70$
7 $80.4$ $80.56$
8 $79.4$ $76.40$
5 Conclusion
This paper introduces LayerCollapse, a novel method for compressing neural networks through regularization. LayerCollapse simplifies a network by merging pairs of linear layers, reducing the network’s depth, the number of parameters, thereby, reducing operational requirements. We introduce a compression aware regularizer that encourages the development of shallower network architectures.
Under rigorous evaluation, LayerCollapse demonstrates state-of-the-art performance in compressing Vision Transformers, improving even improving accuracy in some cases. In comparison to Knowledge Distillation, LayerCollapse consistently demonstrated superior efficiency in both computation and final results, achieving over 10% better performance while utilizing 80% fewer resources. LayerCollapse marks a significant paradigm shift in efficient neural network design and bridges the gap between performance and model size.
Supplementary Material
Appendix A Proof of Equation (4)
Given $Y_{linear}=W_{2}W_{1}X+W_{2}b_{1}+b_{2}$ and $Y_{\alpha}=W_{2}PReLU_{\alpha}(W_{1}X+b_{1})+b_{2}$ we show that $\forall\delta>0$
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (10)
where,
$x^{\delta}\equiv\inf\left\{x\in\mathcal{R}^{n}\mid P\left\{|X|>|x|\right\}<\delta\right\}$ (11)
$C=\sigma_{max}(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2}$
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$.
We define indicator random variable $I$ as:
$I\equiv\begin{cases}1&W_{1}X+b_{1}\geq 0\\ 0&otherwise\end{cases}$ (12)
Writing $Y_{\alpha}$ using (12) we have:
$Y_{\alpha}=IY_{linear}+(1-I)(\alpha W_{2}W_{1}X+\alpha W_{2}b_{1}+b_{2})$ (13)
Using the definition of $Y_{linear}$, the squared distance of $Y_{linear}$ and $Y_{\alpha}$ can be written as:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}=(1-I)^{2}(1-\alpha)^{2}(W_{2}W_{1}X+W_{2}b_{1})^{2}$ (14)
Since $(1-I)^{2}$ is either 00 or $1$ we have:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(W_{2}W_{1}X+W_{2}b_{1})^{2}$ (15)
Using the triangle inequality we can write:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(|W_{2}W_{1}X|^{2}+|W_{2}b_{1}|^{2})$ (16)
It is known that:
$|W_{2}W_{1}X|^{2}\leq\sigma(W_{2}W_{1})^{2}|X|^{2},$ (17)
where, $\sigma(W_{2}W_{1})$ is the largest singular value of the $W_{2}W_{1}$ matrix. Using (17) we have:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(\sigma(W_{2}W_{1})^{2}|X|^{2}+|W_{2}b_{1}|^{2})$ (18)
Using $x^{\delta}$ and $\delta$ as defined in equation (11) we have:
[…]
$(1-\alpha)^{2}(\sigma(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2})\}>1-\delta$
Using equations (19) and (18) with using the $C$ defined in (11) we have:
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (20)
Appendix B Experiment Details
The experimental files are accessible at [https://github.com/sohail2810/LayerCollapse].
B.1 Stand-Alone Compression
Dataset: We utilized the full ImageNet-1k dataset, with image dimensions of $224\times 224$. Training images were augmented with random cropping, horizontal flipping, and rotation (up to 15 degrees). Validation images were not augmented. Normalization factors were model-specific: standard and mean values were set to $0.5$ for all color channels in most models, except for the MLP-Mixer, which used the default ImageNet normalization (mean = ($0.482,0.456,0.406$), std = ($0.229,0.224,0.225$)). The batch size was fixed at 256 for all experiments.
Training Setup: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques. The loss function combined cross-entropy and KL Divergence terms:
$loss=\frac{1}{2}CrossEntropy(p,\hat{p})+\frac{1}{2}$ $KLDivergance(p,\tilde{p})$ (21)
$+Regularization$
where $\hat{p}$ and $\tilde{p}$ represent true and original model’s predicted labels, respectively. Regularization is as defined in Eq. (9).
Layer collapsing was performed sequentially, starting from the last layer. Each layer’s MLP was replaced with our CollapsibleMLP, featuring PReLU activations, and finetuned using Eq. (21) for up to 10 epochs. Following this, we applied the collapse procedure (Algorithm 1). A regularization strength of $0.05$ and an SGD optimizer (momentum $0.9$, initial learning rate $0.0005$) were used. After 5 epochs, the learning rate was reduced to $10^{-4}$.
B.2 LayerCollapse vs. Knowledge Distillation
For both LayerCollapse and Knowledge Distillation, the target architectures were identical. The LayerCollapse setup followed the procedure outlined in Section B.1. The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau. Experiments concluded when no further progress was achievable.
B.3 Ablation Study of Regularization
To ensure comparability, all models, including those without LayerCollapse, shared the same training settings. The SGD optimizer with a starting learning rate of $0.005$ was used, decreasing by a factor of $0.3$ at epochs 75, 150, and 250. A dropout probability of $0.5$ was applied. In LayerCollapse experiments, $0.4$ fraction of layers were regularized with a strength of $0.05$. Hyper-parameter tuning was not a focus of this study.
B.4 LayerCollapse in NLP
We utilized the base Bert model from Hugging Face for NLP experiments. Finetuning employed the Adam optimizer with a learning rate of $0.00002$.